Fermi National Laboratory

The Integrated Luminosity for the period from 6/20/11 to 6/27/11 was 66.87 pb-1 (inverse picobarns). NuMI reported receiving 4.35E18 protons on target during this same period.

Operation Summary for June 24, 2011 - June 27, 2011

 

Friday day June 24

The day shift began with Operations monitoring the Tevatron (TEV) store 8836 (214.13E30) and the Recycler stash (215.8E10), with the Antiproton Source (Pbar) stacking (17.2E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, the Duty Mechanic reported at 8:07 AM that he had cleaned the TEV sector A3 pond water pump strainer and was beginning work on the B3 strainer. NuMI personnel began 40 minutes of target scans at 10:17 AM. Operators began preparing Pbar for an access so that FESS personnel could work on a exhaust fan blower in the afternoon. At 12:44 PM, operators reported that the TEV aborted due to a LCW trip at CUB; water personnel investigated and think that something was valved into the system and contaminated the system; nobody admitted to doing this. Operations reported that the Pbar fan work had been completed and that Pbar had resumed stacking at 2:07 PM. Operations began sending beam to MTA at 2:35 PM for a short period of time. Operators put the TEV through a dry squeeze at 3:16 PM and checked orbits. Operators reported short downtimes due to Booster and Linac RF stations (BRF1 and LRF2).

During the evening shift, Operations established a proton-only store at 4:09 PM. Operations disabled beam to T-978 at 4:41 PM. Operations put the TEV into shot setup at 4:44 PM and established store 8839 at 5:42 PM with a average initial luminosity (AIL) of 382.84E30.

Saturday June 25

The midnight shift began with Operations monitoring store 8839 (135.82E30) and the stash (194.0E10), with Pbar stacking (19.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, around 5:29 AM, the I- Source power supply tripped off due to a LCW resistivity interlock indication. The H- Source supply had also tripped off on the same indication. Operators contacted a Linac person who suggested they check the water resistor skid; operators found its pump was off and its reservoir tank was empty. Operators filled the tank and reported finding a small leak. Mechanical Support personnel investigated and reported at 7:26 AM that they found and replaced a faulty valve on a DI Bottle .

During the day shift, operators reported at 9:57 AM that they were going to terminate store 8839, but then BRF15 tripped off on a LCW indication, which meant the Booster was down two RF stations and could not provide adequate protons for the next store. A Booster RF person came in and brought BRF15 back online. Operations put the TEV into shot setup at 10:45 AM and established store 8840 at 11:53 AM with an AIL of 383.51E30.

During the evening shift, Operations disabled beam to T-978 at 6:12 PM. At 7:22 PM, LRF3 tripped off and would not reset. Operators called in Linac personnel to investigate. At 8:07 PM, operators reported that LRF4 tripped off and would not reset. At 9:25 PM, Operators reported at vacuum burst in the P150 transfer line. Linac personnel reported at 10:59 PM that they had replaced LRF3’s intermediate power amplifier and were beginning to work on LRF4.

Sunday June 26

The midnight shift began with Linac down for a LRF4 intermediate power amplifier replacement, with Operations monitoring store 8840, the stash, and the stack, and with NuMI and MiniBooNE down due to Linac.

During the midnight shift, Linac personnel brought LRF4 online and operators restored beam to MiniBooNE at 12:38 AM and beam to NuMI and Pbar by 12:44 AM. The Duty Electrician and Mechanic installed a portable air conditioning (AC) unit at the B2 service building since the building’s AC unit wouldn’t operate. Operators reported at 6:54 AM that the LRF4 feed-forward was not functioning; Linac personnel came in and adjusted LRF4’s gradient regulator board.

During the day shift, Operations terminated store 8840 at 8:09 AM, put the TEV into shot setup, and established store 8841 at 9:16 AM with an AIL of 361.46E30. MIRF9 and 12 tripped off of 9:58 AM; they reset. BRF7 and 12 tripped off; BRF7 reset, but operators reported some trouble getting BRF12 to reset. Operators bypassed BRF12 and called in Booster RF personnel.

During the evening shift, BRF11 tripped off at 4:06 PM; it reset. Operations disabled beam to T-978 at 6:06 PM. Booster personnel reported that they were bringing BRF12 back on line at 7:52 PM. BRF1, 4, and 8 tripped off during the shift; they all reset.

Monday June 27

The midnight shift began with Operations monitoring store 8841, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, BRF19 tripped off; it reset. Operations terminated store 8841 at 1:22 AM and then the TEV aborted at 1:40 AM due to a HFU at A4 discharging. Operators had the HFU changed out by 2:26 AM. Operations put the TEV into shot setup at 4:36 AM and established store 8844 at 5:41 AM with an AIL of 397.83E30. Operations established beam to T-978 at 7:07 AM. At 7:54 AM, Linac tripped off on a door interlock dropping. Operators confirmed that the door had dropped and also reported that no one was in the area. Operations informed the Interlock technicians.

The Plans for Monday

The plans for today are to repair the Linac interlocks for door #1 and to then go ahead and conduct the necessary interlock safety system tests for Linac. After the interlock testing gets completed, Operations will resume stacking, stashing, and storing.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/27/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory